The voltage V(in volts) that results if an alternating current I (in amps) flows through a resistor R is given thus: V = F(T) = IRsinωt. What is the maxima of the function?


1
Expert's answer
2022-02-03T09:05:23-0500

Solution:

Given, "V = F(t) = IR\sinωt"

"V'= ωIR\cosωt
\\ V''=-ω^2IR\sinωt"

Put V'=0

"\Rightarrow ωIR\cosωt=0
\\ \Rightarrow \cosωt=0
\\ \Rightarrow ωt=\frac{\pi }{2},\:ωt=\frac{3\pi }{2}; 0\le ωt\le 2\pi
\\ \Rightarrow t=\frac{\pi }{2ω},\:ωt=\frac{3\pi }{2w}"

Put these values of t in V' '.

"V''=-ω^2IR\sin(ω\frac{\pi }{2ω})=-ω^2IR<0"

So, maxima exists.

Now, put "t=\frac{\pi }{2ω}" in V.

"V= IR\sin(ω.\frac{\pi }{2ω})
\\ \Rightarrow V=IR(1)
\\\Rightarrow V=IR"

This is maxima.